### v5

Abstract
Computing the edge expansion of a graph is a famously hard combinatorial problem for which there have been many approximation studies. We present two variants of exact algorithms using semidefinite programming (SDP) to compute this constant for any graph. The first variant uses the SDP relaxation first to reduce the search space considerably. The problem is then transformed into instances of max-cut problems which are solved with an SDP-based state-of-the-art solver. Our second variant to compute the edge expansion uses Dinkelbach’s algorithm for fractional programming. This is, we have to solve a parametrized optimization problem and again we use semidefinite programming to obtain solutions of the parametrized problems. Numerical results demonstrate that with our algorithms one can compute the edge expansion on graphs up to 400 vertices in a routine way, including instances where standard branch-and-cut solvers fail. To the best of our knowledge, these are the first SDP-based solvers for computing the edge expansion of a graph.
Keywords: Edge expansion, Cheeger constant, bisection problems, semidefinite programming, parametric submodular minimization
1 Introduction
Let $G=(V,E)$ be a simple connected graph on $n\geq 3$ vertices and with $m$ edges. A cut $(S,S^{\prime})$, for any $\emptyset\neq S\subset V$ and $S^{\prime}=V\setminus S$, in $G$ is a partition of its vertices, and the cut-set $\partial S$ is the edges between $S$ and $S^{\prime}$. The (unweighted) edge expansion, also called the Cheeger constant or isoperimetric number or sparsest cut, of $G$ is a ratio that measures the relative number of edges across any vertex partition. It is defined as
$h(G)$ $=\min_{S}\left\{\frac{|\partial S|}{\min\{|S|,|S^{\prime}|\}}:\emptyset\neq S\subset V\right\}$
[…]
where $\partial S=\{(i,j)\in E:i\in S,\,j\in S^{\prime}\}$ is the cut-set associated with any vertex subset $S\subset V$, and $S^{\prime}=V\setminus S$. This constant is positive if and only if the graph is connected, and the exact value tells us that the number of edges across any cut in $G$ is at least $h(G)$ times the number of vertices in the smaller partition. A weighted definition of edge expansion called the conductance of a graph, is
[…]
where $\operatorname{vol}(S)=\sum_{v\in S}\operatorname{deg}(v)$, and the second equality is due to $\operatorname{vol}(S)+\operatorname{vol}(S^{\prime})=2m$.
Edge expansions arise in the study of expander graphs, for which there is a rich body of literature with applications in network science, coding theory, cryptography, complexity theory, cf. [36, 19, 12]. A graph with $h(G)\geq c$, for some constant $c>0$, is called a $c$-expander. A graph with $h(G)<1$ is said to have a bottleneck since there are not too many edges across it. A threshold for good expansion properties is having $h(G)\geq 1$, which is desirable in many of the above applications. The famous Mihail-Vazirani conjecture [29, 10] in polyhedral combinatorics claims that the graph (1-skeleton) of any 0/1-polytope has edge expansion at least 1. This has been proven to be true for several combinatorial polytopes [29, 20] and bases-exchange graphs of matroids [1], and a weaker form was established recently for random 0/1-polytopes [25]. Lattice polytopes were constructed in [13] with the property that in every dimension their graphs lie on the threshold of being good expanders (i.e., $h(G)=1$).
Computing the edge expansion is related to the uniform sparsest cut problem which asks for computing a cut in the graph with the smallest sparsity, where sparsity is defined as the ratio of the size of the cut to the product of the sizes of the two partitions,
$\phi(G)$ $=\min_{S}\left\{\frac{|\partial S|}{|S||S^{\prime}|}:\emptyset\neq S\subset V\right\}$
[…]
Since $\nicefrac{{n}}{{2}}\leq|S^{\prime}|\leq n$, it holds that $|S||S^{\prime}|\leq n\cdot|S|\leq 2|S||S^{\prime}|$, and hence $h(G)\leq n\cdot\phi(G)\leq 2h(G)$, which implies that the edge expansion problem is related to the sparsest cut problem up to a constant factor of 2. In particular, any cut $(S,S^{\prime})$ that is $\alpha$-approx for $\phi(G)$ (resp. $h(G)$) is a $2\alpha$-approx for $h(G)$ (resp. $\phi(G)$), because $|\partial S|/|S|\leq n|\partial S|/(|S||S^{\prime}|)\leq\alpha\cdot n\cdot\phi(G)\leq 2\alpha\cdot h(G)$.
There are polynomial reductions between $h(G),h_{\operatorname{vol}}(G)$ and $\phi(G)$ and they are all NP-hard to compute [24], in contrast to the minimum-cut of a graph which can be computed in polynomial time. Hence, almost all of the literature on edge expansion is devoted to finding good theoretical bounds. These are generally associated with the eigenvalues of the Laplacian matrix of the graph and form the basis for the field of spectral graph theory (see the monograph [8]). There have also been many approximation studies on this topic [24, 2, 31, 34], and semidefinite optimization (SDP) has been a popular tool in this regard. The best-known approximation for $\phi(G)$ is the famous $\mathcal{O}(\sqrt{\log n})$ factor by Arora et al. [2] which improved upon the earlier $\mathcal{O}({\log n})$-approximation [24]. The analysis is based on an SDP relaxation with triangle inequalities and uses metric embeddings and concentration of measure results. Meira and Miyazawa [28] developed a branch-and-bound algorithm for computing $\phi(G)$ using SDP relaxations and SDP-based heuristics. Recall that $\phi(G)$ is related to $h(G)$ in the approximate sense (up to a factor 2) but not in the exact sense. To the best of our knowledge, there is no exact solution algorithm for $h(G)$.
Contribution and outline
We adopt mathematical programming approaches for numerical computation of $h(G)$. All our approaches make use of tight bounds obtained via semidefinite programming. The first algorithm works in two phases. In the first phase, we split the problem into subproblems and by computing lower and upper bounds for these subproblems, we can exclude a significant part of the search space. In the second phase, we either solve the remaining subproblems to optimality or until a subproblem can be pruned due to the bounds. For the second phase, we transform each subproblem into an instance of a max-cut problem and compute the maximum cut using an SDP-based solver.
The second algorithm we implement uses the idea of Dinkelbach’s algorithm to solve fractional optimization problems. The main concept of this algorithm is to iteratively solve linearly constrained binary quadratic programs. We solve these problems again by transforming them into instances of max-cut and using an SDP-based solver to compute the maximum cut.
We perform numerical experiments on different types of instances which demonstrate the effectiveness of our approaches. To the best of our knowledge, no other algorithms are capable of computing the edge expansion for graphs with a few hundred vertices.
The rest of the paper is structured as follows. In § 2 we formulate the problem as a mixed-binary quadratic program and present an SDP relaxation. § 3 investigates a related problem, namely the $k$-bisection problem. We introduce lower and upper bounds and describe how the $k$-bisection problem can be solved by transforming it to a max-cut problem. The first algorithm (relying on the $k$-bisection problem) for computing $h(G)$ is introduced in § 4, and another algorithm (following Dinkelbach’s idea) in § 5. The performance of all algorithms is demonstrated in § 6, followed by conclusions in § 7.
Notation
The set of $n\times n$ real symmetric matrices is denoted by ${\mathcal{S}}^{n}$. The positive semidefiniteness condition for $X\in\mathcal{S}^{n}$ is written as $X\succeq 0$. The trace of $X$ is written as $\operatorname{tr}(X)$ and defined as the sum of its diagonal elements. The trace inner product for $X,Y\in{\mathcal{S}}^{n}$ is defined as $\langle X,Y\rangle\;=\;\operatorname{tr}(XY)$ and the operator $\operatorname{diag}(X)$ returns the main diagonal of matrix $X$ as a vector. The vector of all ones is $e$ and the matrix of all ones is $J=ee^{\top}$.
[…]
2 Formulations and SDP relaxations
To write an algebraic optimization formulation for cut problems in graphs, we represent a cut $(S,S^{\prime})$ in $G$ by its incidence vector $\chi^{S}\in\{0,1\}^{n}$ which has $\chi^{S}_{i}=1$ if and only if $i\in S$. The cut function is the size of a cut-set, also called the value of the cut, and is equal to
$\zeta(S)=|\partial S|=\sum_{(i,j)\in E}\big{(}\chi^{S}_{i}-\chi^{S}_{j}\big{)}^{2}=\big{(}\chi^{S}\big{)}^{\top}L\chi^{S}\ .$
Any binary vector $x\in\{0,1\}^{n}$ represents a cut in this graph. Denote the set of all cuts with $S$ containing at least one vertex and at most half of the vertices by
=$\mathcal{F}=\left\{x\in\{0,1\}^{n}:1\leq e^{\top}x\leq\frac{n}{2}\right\}.$
Using the common expression $x^{\top}Lx$ for the cut function, the edge expansion problem is
$\begin{split}h(G)&=\min_{x}\;\Big{\{}\dfrac{x^{\top}Lx}{e^{\top}x}:x\in\mathcal{F}\Big{\}}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ &=\min_{x,y}\;\Big{\{}y:\dfrac{x^{\top}Lx}{e^{\top}x}\leq y,\,x\in\mathcal{F}\Big{\}}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ &=\min_{x,y}\;\big{\{}y:x^{\top}Lx-y\,e^{\top}x\leq 0,\,x\in\mathcal{F}\big{\}}.\end{split}$ (1)
The last formulation is a mixed-binary quadratically constrained problem (MIQCP). Standard branch-and-cut solvers may require a large computation time with these formulations even for instances of small to medium size, as we will report in § 6.
Although the focus of this paper is on computing $h(G)$, let us also mention for the sake of completeness that analogous formulations can be derived for the graph conductance (weighted edge expansion) $h_{\operatorname{vol}}(G)$ that was defined in § 1, by optimizing over the set
$\mathcal{F}_{\operatorname{vol}}=\left\{x\in\{0,1\}^{n}:1\leq d^{\top}x\leq m\right\},$
where $d=\operatorname{diag}(D)$ is the vector formed by the vertex degrees. For example, the same steps as in 1 yields the MIQCP
$h_{\operatorname{vol}}(G)=\min_{x,y}\;\big{\{}y:x^{\top}Lx-y\,d^{\top}x\leq 0,\,x\in\mathcal{F}_{\operatorname{vol}}\big{\}}.$
2.1 Semidefinite relaxations
A well-known lower bound for the edge expansion is the spectral bound. It is based on the second smallest eigenvalue of the Laplacian matrix of the graph, namely $h(G)\geq\nicefrac{{\lambda_{2}(L)}}{{2}}$. One way to derive this bound is by considering the following SDP relaxation
$\begin{array}[]{llllll}h(G)\geq&\min_{\widetilde{X},k}{}&\frac{1}{k}\langle L,\widetilde{X}\rangle&~{}=~{}\min_{X}{}&\langle L,X\rangle&\\[1.00006pt] &\text{s.t.}{}&\operatorname{tr}(\widetilde{X})=k&\phantom{~{}=~{}~{}}\text{s.t.}{}&\operatorname{tr}(X)=1\\[1.00006pt] &&\langle J,\widetilde{X}\rangle=k^{2}&&1\leq\langle J,X\rangle\leq\frac{n}{2}\\ &&1\leq k\leq\frac{n}{2}&&X\succeq 0,\\ &&\widetilde{X}\succeq 0\end{array}$ (2)
where $\tilde{X}$ models $xx^{\top}$ and $k=e^{\top}x$. We add the redundant constraint $\langle J,\widetilde{X}\rangle=k^{2}$ and relax $x\in\{0,1\}^{n}$, $\widetilde{X}=xx^{\top}$ to $\widetilde{X}\succeq 0$ to obtain the above SDP relaxation. To eliminate the variable $k$ in the second, equivalent SDP formulation, we scale $X=\frac{1}{k}\widetilde{X}$.
Proposition 2.1.
The optimal solution of the second SDP in (2) is $\nicefrac{{\lambda_{2}(L)}}{{2}}$.
Proof.
First observe that $\frac{1}{n}I$ is a strictly feasible point of the primal SDP and the optimum value is finite, hence strong duality holds. The dual of the second SDP in (2) is
$\max_{v}\;\Big{\{}v_{1}-\frac{n}{2}v_{2}+v_{3}:W=L-v_{1}I+(v_{2}-v_{3})J\succeq 0,\,v_{2},v_{3}\geq 0\Big{\}}.$
The eigenvalue of $W$ with respect to the eigenvector $e$ is $-v_{1}+n(v_{2}-v_{3})$. The other eigenvalues of $W$ are then $\lambda_{i}(L)-v_{1}$ for $2\leq i\leq n$. Therefore, we can write the dual as
$\max_{v}\;\Big{\{}v_{1}-\frac{n}{2}v_{2}+v_{3}:n(v_{2}-v_{3})\geq v_{1},\ \lambda_{2}(L)\geq v_{1},\ v_{2},v_{3}\geq 0\Big{\}},$
which is a linear program with optimal solution $v_{1}=\lambda_{2}(L)$, $v_{2}=\nicefrac{{\lambda_{2}(L)}}{{n}}$ and $v_{3}=0$ and optimal value $\nicefrac{{\lambda_{2}(L)}}{{2}}$.
To strengthen the SDP relaxation 2 we round down the upper bound to $\lfloor\frac{n}{2}\rfloor$ and add the following facet-inducing inequalities of the boolean quadric polytope [32] for $\widetilde{X}$
$0\leq\widetilde{X}_{ij}$ $\leq\widetilde{X}_{ii}$ (3a)
$\widetilde{X}_{i\ell}+\widetilde{X}_{j\ell}-\widetilde{X}_{ij}$ $\leq\widetilde{X}_{\ell\ell}$ (3b)
$\widetilde{X}_{ii}+\widetilde{X}_{jj}-\widetilde{X}_{ij}$ $\leq 1$ (3c)
$\widetilde{X}_{ii}+\widetilde{X}_{jj}+\widetilde{X}_{\ell\ell}-\widetilde{X}_{ij}-\widetilde{X}_{i\ell}-\widetilde{X}_{j\ell}$ $\leq 1,$ (3d)
resulting in the following valid inequalities for $X$
$0\leq{X}_{ij}$ $\leq{X}_{ii}$ (4a)
${X}_{i\ell}+{X}_{j\ell}-{X}_{ij}$ $\leq{X}_{\ell\ell}$ (4b)
${X}_{ii}+{X}_{jj}-{X}_{ij}$ $\leq 1$ (4c)
${X}_{ii}+{X}_{jj}+{X}_{\ell\ell}-{X}_{ij}-{X}_{i\ell}-{X}_{j\ell}$ $\leq 1$ (4d)
for all $1\leq i,\ j,\ \ell\leq n$. Note, that in 4c and 4d we have to replace $\frac{1}{k}$ in the rhs by its upper bound 1 in order to obtain a formulation without $k$. Therefore, we cannot expect these inequalities to strengthen the SDP relaxation significantly.
2.2 Illustrative examples for motivation
We motivate our algorithm by considering the example of the graph of the grlex polytope, which is described in [13]. Table 1 compares different lower bounds on $h(G)$ for these graphs. The first column indicates the dimension of the polytope and the second column lists the number of vertices in the associated graph. The third column gives the edge expansion that is known to be one for these graphs in all dimensions [13]. The spectral bound is displayed in the fourth column. Column 5 lists the optimal value of the SDP relaxation 2 strengthened by inequalities 4 derived from the boolean quadric polytope. Column 6 displays a lower bound that is very easy to compute: the minimum cut of the graph divided by the largest possible size of the smaller set of the bipartition of the vertices, that is $\lfloor\frac{n}{2}\rfloor$. In the last column, the minimum of the lower bounds $\ell_{k}$ for $1\leq k\leq\lfloor\frac{n}{2}\rfloor$ is listed with $\ell_{k}$ being a bound related to the solution of 2 for $k$ fixed. The definition of $\ell_{k}$ follows in § 3.1.
The numbers in the table show that some of these bounds are very weak, in particular, if the number of vertices increases. Interestingly, if we divide the edge expansion problem into $\lfloor\frac{n}{2}\rfloor$ many subproblems with fixed denominator (as we did to obtain the numbers in column 6) the lower bound we obtain by taking the minimum over all SDP relaxations for the subproblems seems to be stronger than the other lower bounds presented in Table 1. We will, therefore, take this direction of computing the edge expansion, namely, we will compute upper and lower bounds on the problem with fixed $k$. Using these bounds will allow to exclude a (hopefully) large number of potential sizes $k$ of the smaller partition. This will leave us with computing the maximum cut of a graph with fixed sizes of the partition $k$ and $n-k$ for a few values of $k$ only.
$d$
$n$
$h(G)$
$\nicefrac{{\lambda_{2}}}{{2}}$
2 & 4
$\zeta_{\min}(G)/\lfloor\frac{n}{2}\rfloor$
$\min_{k}\,\ell_{k}$
2 4 1 1.0000 1.0000 1.0000 1.0000
3 7 1 0.7929 0.8750 1.0000 1.0000
4 11 1 0.6662 0.7857 0.8000 1.0000
5 16 1 0.5811 0.7273 0.6250 1.0000
6 22 1 0.5231 0.6875 0.5455 1.0000
7 29 1 0.4820 0.6591 0.5000 1.0000
8 37 1 0.4516 0.6379 0.4444 1.0000
Table 1: Comparison of lower bounds for graphs from the grlex polytope in dimension $d$.
3 Fixing the size $k$: Bisection problem
If the size $k$ of the smaller set of the partition of an optimum cut is known, the edge expansion problem would result in a scaled bisection problem. That is, we ask for a partition of the vertices into two parts, one of size $k$ and one of size $n-k$, such that the number of edges joining these two sets is minimized. This problem is NP-hard [11] and has the following formulations for any $k\in\{1,2,\dots,\lfloor\frac{n}{2}\rfloor\}$,
$\begin{array}[]{rll}h_{k}=\frac{1}{k}&\min_{x}&x^{\top}Lx\\[1.99997pt] &\text{s.t.}&e^{\top}x=k\\ &&x\in\{0,1\}^{n},\end{array}$ (5)
but standard branch-and-cut solvers can solve these in reasonable time only for small-sized graphs.
Since SDP-based bounds have been shown to be very strong for partitioning problems, cf. [27, 38, 21, 39], we exploit these bounds with our first algorithm to compute the edge expansion. In the subsequent sections, we describe how to obtain lower and upper bounds on $h_{k}$ (§ 3.1 and 3.2). We then present in § 3.3, how to transform the bisection problem into an instance of a max-cut problem which is then solved using the state-of-the-art solver BiqBin [16]. For completeness, a description of BiqBin is given in Appendix A.
3.1 SDP lower bounds for the bisection problem
After squaring the linear equality constraint in problem 5 and employing standard lifting and relaxation techniques, we obtain the following SDP relaxation that is generally computationally cheap to solve,
$\begin{array}[]{rll}\ell_{\text{bisect}}(k)=&\min_{X,x}&\langle L,X\rangle\\[1.99997pt] &\text{s.t.}&\operatorname{tr}(X)=k\\ &&\langle J,X\rangle=k^{2}\\ &&\operatorname{diag}(X)=x\\ &&\begin{pmatrix}1&&x^{\top}\\ x&&X\end{pmatrix}\succeq 0.\end{array}$ (6)
Since the bisection for a given simple unweighted graph has to be an integer, we get the following lower bound on the scaled bisection $h_{k}$,
$h_{k}\geq\ell_{k}=\frac{\lceil\ell_{\text{bisect}}(k)\rceil}{k}.$ (7)
There are several ways to strengthen the above relaxation of the bisection problem. In [39] a vector lifting SDP relaxation, tightened by non-negativity constraints, has been introduced. In our setting, this results in the following doubly non-negative programming (DNN) problem,
$\begin{split}\min_{X}~{}~{}&\langle L,X^{11}+X^{22}\rangle\\ \textrm{s.t.}~{}~{}&\operatorname{tr}(X^{11})=k,~{}\langle J,X^{11}\rangle=k^{2}\\ &\operatorname{tr}(X^{22})=n-k,~{}\langle J,X^{22}\rangle=(n-k)^{2}\\ &\operatorname{diag}(X^{12})=0,~{}\operatorname{diag}(X^{21})=0,~{}\langle J,X^{12}+X^{21}\rangle=2k(n-k)\\ &X=\begin{pmatrix}1&(x^{1})^{\top}&(x^{2})^{\top}\\ x^{1}&X^{11}&X^{12}\\ x^{2}&X^{21}&X^{22}\\ \end{pmatrix}\succeq{0},~{}~{}x^{i}=\operatorname{diag}(X^{ii}),~{}~{}i=1,2\\ &X\geq 0,\end{split}$ (8)
where $X$ is a matrix of size $(2n+1)\times(2n+1)$. This relaxation can be further strengthened by cutting planes from the Boolean Quadric Polytope. In particular, we want to add the inequalities
$X_{i\ell}+X_{j\ell}\leq X_{\ell\ell}+X_{ij}$ (9)
as Meijer et al. [27] demonstrated that these inequalities are the most promising ones to improve the bound.
The DNN relaxation 8 cannot be solved by standard methods due to the large number of constraints. The additional cutting-planes 9 make the SDP relaxation extremely difficult to solve already for medium-sized instances. Meijer et al. [27] apply facial reduction to the SDP relaxation which leads to a natural way of splitting the set of variables into two blocks. Using an alternating direction method of multipliers (ADMM) provides (approximate) solutions to this relaxation even for graphs with up to 1000 vertices. The steps to be performed in this ADMM algorithm result in projections onto the respective feasible sets. For projections onto polyhedra, Dykstra’s projection algorithm is used. A careful selection of non-overlapping cuts, warm starts, and an intelligent separation routine are further ingredients of this algorithm in order to obtain an efficient solver for the SDP 8 enhanced with inequalities 9. A post-processing algorithm is also introduced to guarantee a valid lower bound. Using this algorithm, we can compute strong lower bounds for each $k$ with reasonable computational effort.
3.2 A heuristic for the bisection problem
The graph bisection problem can also be written as a quadratic assignment problem (QAP) [22]. To do so, we set the weight matrix $W$ to be the Laplacian matrix $L$ of the graph and the distance matrix $\widetilde{D}$ to be a matrix with a top left block of size $k$ with all ones and the rest zero. The resulting QAP for this weight and distance matrix is
$\min_{\pi\in\Pi_{n}}\ \sum_{i=1}^{n}\sum_{j=1}^{n}W_{i,j}\widetilde{D}_{\pi(i),\pi(j)}=\min_{\pi\in\Pi_{n}}\ \sum_{i=1}^{k}\sum_{j=1}^{k}L_{\pi^{-1}(i),\pi^{-1}(j)}=kh_{k}.$
In this formulation, the vertices mapped to values between 1 and $k$ by the permutation $\pi$ are chosen to be in the set of size $k$ in the partition. To compute an upper bound $u_{k}$ on $h_{k}$, we can use any heuristic for the QAP and divide the solution by $k$.
Simulated annealing is a well-known heuristic to compute an upper bound for the QAP, we implement the algorithm introduced in [7]. We use a slightly different parameter setting that we determined via numerical experiments. That is, the initial temperature is set to $\nicefrac{{k^{2}}}{{\binom{n}{2}}}\cdot\operatorname{tr}(L)$, the number of transformation trials at constant temperature is initially set to $n$ and increased by a factor of 1.15 after each cycle, and the cooling factor is set to 0.7. After every trial, we additionally perform a local search strategy to find the local minimum.
Other well-performing heuristics for the QAP are tabu search, genetic algorithms, or algorithms based on the solution of the SDP relaxation like the hyperplane rounding algorithm. Some of these heuristics have the potential to be superior to simulated annealing. However, as we will see later in the study of our experiments, the bounds we obtain with the simulated annealing heuristic are almost always optimal for the size of instances we are interested in.
3.3 Transformation to a max-cut problem
One variant to solve the $k$-bisection problem is to implement a branch-and-bound algorithm with the aforementioned lower and upper bounds as presented in [14]. We are going to use a different approach to solving the graph bisection problem. Namely, we transform it to a max-cut problem and then take advantage of using a well-established and performant max-cut solver, e.g. the open source parallel solver BiqBin [16], see also Appendix A. To do so, we first need to transform the bisection problem into a quadratic unconstrained binary problem (QUBO).
Lemma 3.1.
Let $\tilde{x}\in\{0,1\}^{n}$ such that $e^{\top}\tilde{x}=k$, and choose $\mu_{k}$ such that $\mu_{k}>\tilde{x}^{\top}L\tilde{x}$. Then
$h_{k}=\frac{1}{k}\min_{x}\;\big{\{}x^{\top}(L+\mu_{k}J)x-2\mu_{k}k\,e^{\top}x+\mu_{k}k^{2}:x\in\{0,1\}^{n}\big{\}}.$
Proof.
First note that
$x^{\top}(L+\mu_{k}ee^{\top})x-2\mu_{k}k\,e^{\top}x+\mu_{k}k^{2}=x^{\top}Lx+\mu_{k}\|e^{\top}x-k\|^{2}.$
Let $x\in\{0,1\}^{n}$. Then we have
[…]
Note that $e^{\top}x-k$ is integer for $x\in\{0,1\}^{n}$. Hence, for any infeasible $x\in\{0,1\}^{n}$, the objective is greater than the given upper bound $\tilde{x}^{\top}L\tilde{x}$, and therefore the minimum can only be attained for $x\in\{0,1\}^{n}$ with $e^{\top}x=k$. ∎
Barahona et al. [5] showed that any QUBO problem can be reduced to a max-cut problem by adding one additional binary variable. In our context, this means the following.
Corollary 3.2.
Let $G=(V,E)$ and let $G^{\prime}$ be the complete graph with vertex set $V\cup\{v_{0}\}$. Let the weights $c_{uw}$ on the edges of $G^{\prime}$ be as follows.
$c_{uw}=\begin{cases}4\mu_{k}(n-2k)&\text{ if }u\in V(G)\text{ and }w=v_{0}\\ 4\mu_{k}-1&\text{ if }uw\in E(G)\\ 4\mu_{k}&\text{ if }uw\not\in E(G)\end{cases}$
Then the minimum bisection of $G$ where one side of the cut has size $k$ is equal to $\mathtt{offset}-\operatorname{max-cut}(G^{\prime})$, where $\mathtt{offset}=4\mu_{k}(n-k)^{2}$.
Since max-cut solvers can benefit from edge weights of the input graph being integer, a possible choice for $\mu_{k}$ is an upper bound on the bisection problem plus $\nicefrac{{1}}{{4}}$. Note that we choose $\mu_{k}$ to be as small as possible by doing so.
The formulation of the max-cut problem in ±$\pm 1$ variables additionally requires $x_{v}=1$ to hold. Because of the symmetry of the cut, we can omit this constraint. Due to our choice of the penalty parameter, it holds that on one side of the maximum cut, there are exactly $k+1$ vertices, including vertex $v$. These $k$ vertices on the same side as $v$ are the vertices in the set of size $k$ in the optimum of the bisection problem.
To summarize, we can solve the bisection problem by solving a dense max-cut problem. With this, we now have all the tools needed for our new split & bound approach.
4 Split & bound
We now assemble the tools developed in § 3 to compute the edge expansion of a graph by splitting the problem into $\lfloor\frac{n}{2}\rfloor$ many bisection problems. Since the bisection problem is NP-hard as well, we want to reduce the number of bisection problems we have to solve exactly as much as possible. To do so, we start with a pre-elimination of the bisection problems. This procedure aims to exclude subproblems unnecessary for the computation of the edge expansion of the graph. Computing the edge expansion by considering the remaining values of $k$ is summarized in Algorithm 2 below. We now explain the pre-elimination step and further ingredients of our algorithm.
4.1 Pre-elimination
The size $k$ of the smaller set of the partition can theoretically be any value from $1$ to $\lfloor\frac{n}{2}\rfloor$. However, it can be expected that for some candidates, one can quickly check that the optimal solution cannot be attained for that $k$. As a first quick check, we use the cheap lower bound $\ell_{k}$ obtained by solving the SDP 6 in combination with the upper bound introduced in the § 3.2. We do not need to further consider values of $k$ where the lower bound $\ell_{k}$ of the scaled bisection problem is already above an upper bound on the edge expansion. A pseudo-code of this pre-elimination step is given in Algorithm 1.
1 for $k\in\{1,\dots,\lfloor\frac{n}{2}\rfloor\}$ do
2       Compute an upper bound $u_{k}$ using the heuristic from § 3.2;
3       Compute the lower bound $\ell_{k}$ from 7 by solving the cheap SDP 6;
4Global upper bound $u^{*}\coloneqq\min\big{\{}u_{k}\colon 1\leq k\leq\lfloor\frac{n}{2}\rfloor\big{\}}$;
5 if $\min_{k}\ell_{k}=u^{*}$ then
6      $\mathcal{I}=\emptyset$, $h(G)=u^{*}$;
7else
8       $\mathcal{I}\coloneqq\big{\{}k\in\{1,\dots,\lfloor\frac{n}{2}\rfloor\}\colon\ell_{k}<u^{*}\big{\}}$;
return $\mathcal{I}$, $u_{k}$ for $k\in\mathcal{I}$, $u^{*}$
Algorithm 1 Pre-eliminate certain values of $k$
The hope is that many values of $k$ can be excluded from computing the edge expansion. Clearly, this heavily depends on the instance itself, as in the worst case, it might happen that for many different values of $k$, the value of $h_{k}$ is close to the optimum.
We can further reduce the number of candidates for $k$ by computing a tighter lower bound ~$\tilde{\ell}_{k}$ by solving the DNN relaxation 8 with additional cutting planes. In our implementation we do not compute ~$\tilde{\ell}_{k}$ as part of the pre-elimination but use the lower bound obtained from the solution in the root node of the max-cut solver.
Impact of pre-elimination on sample instances
Figures 1 and 2 display the bounds associated with four different graphs. For the graph of the grevlex polytope in dimension 7, considering the bounds $u_{k}$ and ~$\tilde{\ell}_{k}$ the only candidates for $k$ where the optimal solution can be found are 12 and 14. For the grevlex polytope in dimension 8, the sizes 17 and 18 remain as the only candidates. Also for a graph associated to a randomly generated 0/1-polytope and to a network graph, about 2/3 of the potential values of $k$ can be excluded already by considering the cheap lower bound $\ell_{k}$.
Figure 1: Lower and upper bounds for each $k$.
Figure 2: Lower and upper bounds for each $k$.
4.2 Stopping exact computations early and updating $u^{*}$
All values of $k$ that are not excluded in the pre-elimination step have to be further examined. For those values we continue to run the max-cut solver to compute the scaled bisection $h_{k}$. We can stop the branch-and-bound algorithm as soon as the lower bound (on the scaled $k$-bisection problem) of all open nodes in the branch-and-bound tree is larger or equal to $u^{*}$. (Remember that $u^{*}$ is an upper bound on the edge expansion of the graph but not necessarily an upper bound on $h_{k}$.) A simple way to implement this stopping criterion is to initialize the algorithm for the $k$-bisection problem with $\lceil u^{*}k\rceil$ as an “artificial” upper bound. For the max-cut solver, this translates to an “artificial” lower bound of $\mathtt{offset}-\lceil u^{*}k\rceil$.
In case $h_{k}<u^{*}$, we can update $u^{*}$ which might lead to eliminating further values from $\mathcal{I}$. This fact is also part of the motivation for the order of choosing $k$ for computing $h_{k}$, as described in the next section.
4.3 Order of selecting values $k$ from $\mathcal{I}$
We consider the order of computing $h_{k}$ in ascending order based on their upper bounds $u_{k}$. The motivation for this choice is as follows.
Remember that $u^{*}=\min_{k}u_{k}$ is a global upper bound on the edge expansion. The most promising values for $k$ to even further improve this bound are those with small $u_{k}$. Therefore, before starting the max-cut solver for the values $k$ left after pre-elimination, we do another 30 trials of simulated annealing for each of these to hopefully further improve the upper bound.
Moreover, we run the exact computation of $h_{k}$ in this order since also during the branch-and-bound algorithm, the upper bound $u_{k}$ might drop further and this will improve the global upper bound $u^{*}$ most likely for those candidates with small $u_{k}$.
An improvement of the upper bound $u^{*}$ means that there is a possibility to further eliminate values $k$ from $\mathcal{I}$. But even for values $k$ that can not be eliminated, we obtain smaller artificial upper bounds and hence the computation of these bisection problems may be stopped earlier.
We summarize all the steps in Algorithm 2.
1 $\mathcal{I}$, $u_{k}$ for $k\in\mathcal{I}$, $u^{*}$leftarrow pre-elimination Algorithm 1;
2 for $k\in\mathcal{I}$ do
3       Run heuristic from § 3.2 and update $u_{k}$;
4       if $u_{k}<u^{*}$ then
5            $u^{*}\leftarrow u_{k}$;
6             update $\mathcal{I}$;
7
8for $k\in\mathcal{I}$, consider $k$ in ascending order of $u_{k}$ do
9       transform the instance to a max-cut instance;
10       initialize the lower bound for max-cut as $\mathtt{offset}-\lceil u^{*}k\rceil$;
11       compute $h_{k}$ using the max-cut solver;
12       if $h_{k}<u^{*}$ then
13            $u^{*}\leftarrow h_{k}$;
14             update $\mathcal{I}$
15
[…]
Algorithm 2 Split & bound
4.4 Algorithmic verification of lower bound
We close this section by addressing the important consideration that we are not interested in the exact value of the edge expansion in some applications, but want to check whether certain values are valid lower bounds on $h(G)$. A lower bound $c\leq h(G)$, for some constant $c>0$, means that the graph is a $c$-expander. The value of this lower bound means that the graph expands by at least that much. This also arises in the context of the Mihail-Vazirani conjecture on 0/1-polytopes where one wants to check whether $h(G)\geq 1$ where $G$ is the graph of a 0/1-polytope.
Our split & bound algorithm can also be used to verify a lower bound.
Proposition 4.1.
Let $\upsilon$ be a given scalar and suppose we initialise Algorithm 2 with $u^{\ast}=\upsilon$. Then $\upsilon$ is a valid lower bound on $h(G)$ if and only if the algorithm terminates without updating $u^{\ast}$.
Proof.
Assume we initialize $u^{\ast}=\upsilon$. If we find a better upper bound (or some computed value $h_{k}$ is smaller than $\upsilon$), this is a certificate that the given value $\upsilon$ is not a valid lower bound since we found a better solution. Otherwise, if the upper bound never gets updated, this means the provided bound is indeed a valid lower bound on the edge expansion of the graph. ∎
5 Parametric optimization
Another approach to compute $h(G)$ is following a discrete Newton-Dinkelbach algorithm. Dinkelbach [9] gave a general classical framework to solve (non)-linear fractional programs. The program one aims to solve is $\min_{x\in\mathcal{F}}\,f(x)$, where the objective $f$ is a fraction of (non)-linear functions. In our case this is
$f(x)=\frac{x^{\top}Lx}{e^{\top}x},\quad\text{and}\quad\mathcal{F}=\Big{\{}x\in\{0,1\}^{n}:1\leq e^{\top}x\leq\frac{n}{2}\Big{\}}.$ (10)
The main component of the algorithm is to form the following parametrized objective function
$g_{\gamma}(x)=x^{\top}Lx-\gamma e^{\top}x,$
and the corresponding parametrized optimization problem
$P(\gamma)=\min_{x}\;\{g_{\gamma}(x):x\in\mathcal{F}\},\qquad\gamma\geq 0.$
This problem then has the following useful properties.
Proposition 5.1.
$P(0)=\zeta_{\min}(G)$ and $P$ is a strictly decreasing concave piecewise linear function over + whose unique root is equal to $h(G)$. Consequently,
$h(G)=\max_{\gamma}\big{\{}\gamma\colon g_{\gamma}(x)\geq 0\big{\}}=\min_{\gamma}\big{\{}\gamma\colon g_{\gamma}(x)\leq 0\big{\}}.$
Proof.
We have
$P(0)$ $=\min_{x}\big{\{}x^{\top}Lx:x\in\mathcal{F}\big{\}}$
$=\min_{S}\big{\{}\lvert{\partial S}\rvert:\emptyset\neq S\subset V,|S|\leq\nicefrac{{n}}{{2}}\big{\}}$
$=\min_{S}\big{\{}\lvert{\partial S}\rvert:\emptyset\neq S\subset V\big{\}}$
$=\zeta_{\min}(G)$
[…]
This implies that the edge expansion of a graph can be computed using a root-finding algorithm for the function $P$. One evaluation of $P$ for a given $\gamma$ still means solving a binary quadratic problem with two linear inequalities. Hence, reducing this number of evaluations is crucial to compute $h(G)$ in reasonable time. There are several strategies to do so, such as binary search. Our approach is to evaluate $P$ starting with $\gamma_{1}$ equal to some good upper bound on $h(G)$ (in our experiments, we used our heuristic from § 3.2). We are already done if we have found the optimum with our heuristic, that is when $P(\gamma_{1})=0$. Otherwise, there is some $x^{1}\in\mathcal{F}$ such that $g_{\gamma}(x^{1})<0$ and therefore $f(x^{1})<\gamma$. This means that $f(x^{1})$ is a better upper bound than $\gamma_{1}$. Hence, we now set $\gamma_{2}=f(x^{1})$ and repeat until we find the optimum as described in Algorithm 3. Since $P(\gamma)<0$ if and only if $\gamma>h(G)$, the stopping criterion is checking whether $P(\gamma)<0$ at the current iterate.
Input: graph $G$, upper bound $\gamma_{1}\geq h(G)$ from heuristic
Output: edge expansion $h(G)$
1 $i=1$;
2 while $P(\gamma_{i})<0$ do
3       $x_{i}\in\arg\min_{x\in\mathcal{F}}g_{\gamma_{i}}(x)$;
4       $\gamma_{i+1}=f(x_{i})$;
5       $i=i+1$;
[…]
Algorithm 3 Discrete Newton-Dinkelbach algorithm for edge expansion
The superlinear convergence rate of Dinkelbach’s algorithm was established in [37]. We derive a similar convergence result for Algorithm 3.
Theorem 5.2.
Algorithm 3 terminates with the optimal value after finitely many steps, the rate of convergence is superlinear.
To prove the convergence rate of Theorem 5.2, we first need the following two lemmas.
Lemma 5.3.
Let $\gamma^{\prime}$, $\gamma^{\prime\prime}\in$ and $x^{\prime}$, $x^{\prime\prime}\in\mathcal{F}$ be the optimal solutions of $P(\gamma^{\prime})$ and $P(\gamma^{\prime\prime})$, then
$f(x^{\prime})-f(x^{\prime\prime})\leq\big{(}P(\gamma^{\prime\prime})-(\gamma^{\prime}-\gamma^{\prime\prime})e^{\top}x^{\prime\prime}\big{)}\Big{(}\frac{1}{e^{\top}x^{\prime}}-\frac{1}{e^{\top}x^{\prime\prime}}\Big{)}.$
Proof.
By the optimality of $x^{\prime}$ for $P(\gamma^{\prime})$ it holds that
$x^{\prime\top}Lx^{\prime}-\gamma^{\prime}e^{\top}x^{\prime}\leq x^{\prime\prime\top}Lx^{\prime\prime}-\gamma^{\prime}e^{\top}x^{\prime\prime}.$
Dividing both sides by $e^{\top}x^{\prime\prime}$ and rearranging yields
$f(x^{\prime})\leq\frac{x^{\prime\prime\top}Lx^{\prime\prime}}{e^{\top}x^{\prime}}+\gamma^{\prime}\Big{(}1-\frac{e^{\top}x^{\prime\prime}}{e^{\top}x^{\prime}}\Big{)}.$
Hence, we get that
$f(x^{\prime})-f(x^{\prime\prime})$ $\leq\frac{x^{\prime\prime\top}Lx^{\prime\prime}}{e^{\top}x^{\prime}}+\gamma^{\prime}\Big{(}1-\frac{e^{\top}x^{\prime\prime}}{e^{\top}x^{\prime}}\Big{)}-\frac{x^{\prime\prime\top}Lx^{\prime\prime}}{e^{\top}x^{\prime\prime}}$
$=(x^{\prime\prime\top}Lx^{\prime\prime}-\gamma^{\prime}e^{\top}x^{\prime\prime})\Big{(}\frac{1}{e^{\top}x^{\prime}}-\frac{1}{e^{\top}x^{\prime\prime}}\Big{)}$
$=(x^{\prime\prime\top}Lx^{\prime\prime}-\gamma^{\prime\prime}e^{\top}x^{\prime\prime}+\gamma^{\prime\prime}e^{\top}x^{\prime\prime}-\gamma^{\prime}e^{\top}x^{\prime\prime})\Big{(}\frac{1}{e^{\top}x^{\prime}}-\frac{1}{e^{\top}x^{\prime\prime}}\Big{)}$
$=\big{(}P(\gamma^{\prime\prime})-(\gamma^{\prime}-\gamma^{\prime\prime})e^{\top}x^{\prime\prime}\big{)}\Big{(}\frac{1}{e^{\top}x^{\prime}}-\frac{1}{e^{\top}x^{\prime\prime}}\Big{)}.$
Lemma 5.4.
Let $x^{\prime}$ and $x^{\prime\prime}$ be optimal solutions of $P(\gamma^{\prime})$ and $P(\gamma^{\prime\prime})$, then for $\gamma^{\prime\prime}<\gamma^{\prime}$ it holds that $e^{\top}x^{\prime\prime}\leq e^{\top}x^{\prime}$.
Proof.
From the optimality of $x^{\prime}$ and $x^{\prime\prime}$ it follows that
$x^{\prime\top}Lx^{\prime}-\gamma^{\prime}e^{\top}x^{\prime}$ $\leq x^{\prime\prime\top}Lx^{\prime\prime}-\gamma^{\prime}e^{\top}x^{\prime\prime}\text{\quad and}$
$x^{\prime\prime\top}Lx^{\prime\prime}-\gamma^{\prime\prime}e^{\top}x^{\prime\prime}$ $\leq x^{\prime\top}Lx^{\prime}-\gamma^{\prime\prime}e^{\top}x^{\prime}.$
Adding the above two inequalities yields
$(\gamma^{\prime\prime}-\gamma^{\prime})e^{\top}x^{\prime}\leq(\gamma^{\prime\prime}-\gamma^{\prime})e^{\top}x^{\prime\prime}$
and hence the above claim holds. ∎
Proof of Theorem 5.2.
Let $x^{*}\in\mathcal{F}$ be the optimum of the edge expansion problem, i.e., $f(x^{*})=h(G)$ and let $\gamma^{*}=f(x^{*})$. From Proposition 5.1 we know that $P$ is a strictly decreasing piecewise linear function and therefore the algorithm terminates after finitely many iterations with value $\gamma^{*}$.
Let further $\gamma_{i}$ be the upper bound on $h(G)$ to check in the $i$-th iteration of Algorithm 3. From Lemma 5.3, we get that
$\gamma_{i+1}-\gamma^{*}\leq(\gamma_{i}-\gamma^{*})\Big{(}1-\frac{e^{\top}x^{*}}{e^{\top}x_{i}}\Big{)}$
holds, since $P(\gamma^{*})=0$. The sequence
$\Big{(}1-\frac{e^{\top}x^{*}}{e^{\top}x_{i}}\Big{)}$
is strictly decreasing (and converging to 0) as proved in Lemma 5.4. Therefore, the convergence rate of Algorithm 3 is superlinear. ∎
5.1 Solving the parametrized optimization problem
Evaluating $P(\gamma)$ requires solving a binary quadratic problem with two linear inequality constraints which is in general NP-hard.
Most solvers for binary quadratic programs benefit from input data given as integer as this aids the performance of the underlying branch-and-bound algorithm. Since we only consider rational values for $\gamma$, we introduce the following parametric optimization problem
$Q(\gamma)=\min_{x}\{\gamma_{d}x^{\top}Lx-\gamma_{n}e^{\top}x:x\in\mathcal{F}\}$
for $\gamma=\nicefrac{{\gamma_{n}}}{{\gamma_{d}}}$ with integers $\gamma_{n}\geq 0$ and $\gamma_{d}>0$. Observe that $Q(\gamma)=\gamma_{d}P(\gamma)$ and all considerations from above apply to this new formulation as well.
Two performant parallel state-of-the-art solvers for binary quadratic programs with linear equality constraints are BiqBin [16] and BiqCrunch [23]. BiqBin first transforms the problem into a QUBO in a pre-processing phase and then solves the equivalent max-cut problem with its max-cut solver. To solve $Q(\gamma)$, we make use of the problem specific properties and directly transform it into a max-cut problem ourselves. The first step towards achieving this is to obtain an exact formulation as a QUBO using binary encoding of the slack variables and the penalty parameter suggested in [17, Thm. 15]. To aid our derivation, let us denote the integer $n_{s}$ and vector $v^{n_{s}}\in^{n_{s}+1}$ by
$n_{s}=\Big{\lceil}\log_{2}\Big{\lfloor}\frac{n}{2}\Big{\rfloor}\Big{\rceil}-1,\qquad v^{n_{s}}_{i}=2^{i-1}\text{ for all }i.$
Proposition 5.5.
Let =$\mathcal{F}=\{x\in\{0,1\}^{n}:1\leq e^{\top}x\leq\frac{n}{2}\}$ and $v^{n_{s}}\in^{n_{s}+1}$ with $v^{n_{s}}_{i}=2^{i-1}$ and $n_{s}=\lceil\log_{2}(\lfloor\frac{n}{2}\rfloor)\rceil-1$. Then
=$\mathcal{F}=\Big{\{}x\in\{0,1\}^{n}:e^{\top}x-\alpha^{\top}v^{n_{s}}=1,\ e^{\top}x+\beta^{\top}v^{n_{s}}=\Big{\lfloor}\frac{n}{2}\Big{\rfloor},\ \alpha,\beta\in\{0,1\}^{n_{s}+1}\Big{\}}.$
Proof.
For any $x\in\mathcal{F}$ it holds that $e^{\top}x=1+s=\lfloor\frac{n}{2}\rfloor-t$ for some slack variables $s$ and $t$ with $10\leq s,t\leq\lfloor\frac{n}{2}\rfloor-$ 1. In fact, any upper bound on $s$ and $t$ greater or equal than $\lfloor\frac{n}{2}\rfloor-1$ is fine, since from $e^{\top}x=1+s$ and $s\geq 0$ it follows that $e^{\top}x\geq 1$ and from $e^{\top}x=\lfloor\frac{n}{2}\rfloor-t$ and $t\geq 0$ it follows that $e^{\top}x\leq\lfloor\frac{n}{2}\rfloor$. The smallest possible value for $n_{s}$ is $\lceil\log_{2}(\lfloor\frac{n}{2}\rfloor)\rceil-1$, since this gives an upper bound of $12^{n_{s}+1}-$ 1 on $s$ and $t$. ∎
Proposition 5.6.
Let $x^{\prime}\in\mathcal{F}$ with $\frac{x^{\prime\top}Lx^{\prime}}{e^{\top}x^{\prime}}=\frac{\gamma_{n}}{\gamma_{d}}=\gamma$ and $\gamma_{d}>0$. The problem $Q(\gamma)$ can then be equivalently formulated as the following QUBO,
[…]
with $\sigma>\gamma_{n}n$.
Proof.
Let $g(x,\alpha,\beta)$ denote the objective function of (11). For any feasible vector $x\in\mathcal{F}$ there exist uniquely defined $\alpha_{x}$, $\beta_{x}$ such that $g(x,\alpha_{x},\beta_{x})=\gamma_{d}x^{\top}Lx-\gamma_{n}e^{\top}x$. Thus, the objective function of $Q(\gamma)$ and (11) coincide for $x\in\mathcal{F}$. Moreover, for $x^{\prime}$ it holds that $g(x^{\prime},\alpha_{x^{\prime}},\beta_{x^{\prime}})=0$.
For $x\in\{0,1\}^{n}\setminus\mathcal{F}$ there do not exist $\alpha$, $\beta\in\{0,1\}^{n_{s}+1}$ such that both equalities $e^{\top}x-\alpha^{\top}v^{n_{s}}=1$ and $e^{\top}x+\beta^{\top}v^{n_{s}}=\big{\lfloor}\frac{n}{2}\big{\rfloor}$ are satisfied, as one of the slack variables has to be negative in order to fulfill the constraints. Additionally, since $L$ is positive semidefinite we can conclude that
$g(x,\alpha,\beta)\geq-\gamma_{n}n+\sigma>0=g(x^{\prime},\alpha_{x^{\prime}},\beta_{x^{\prime}}).$
Therefore, $Q(\gamma)$ is an equivalent formulation of (11). ∎
The unconstrained binary quadratic program (11) can again be transformed to a max-cut problem, as explained in [5] for example. Applied to our problem we obtain the following result.
Corollary 5.7.
Let $G=(V,E)$ and let $G^{\prime\prime}$ be the graph with vertices from $V$ plus the vertices $v_{0},v_{\alpha_{0}},\dots,v_{\alpha_{n_{s}}},v_{\beta_{0}},\dots,v_{\beta_{n_{s}}}$ for the variable vectors $\alpha$ and $\beta$.
Let the weights $c_{uw}$ on the edges of $G^{\prime\prime}$ be as follows.
$c_{uv_{0}}=\begin{cases}2\sigma(n-1-\lfloor\frac{n}{2}\rfloor)-\gamma_{n}&\text{ if }u\in V(G)\\ 2\sigma(2^{n_{s}}-\frac{n-1}{2})2^{i}&\text{ if }u=v_{\alpha_{i}}\\ 2\sigma(2^{n_{s}}-\lfloor\frac{n}{2}\rfloor+\frac{n-1}{2})2^{i}&\text{ if }u=v_{\beta_{i}}\end{cases}$
$c_{uw}=\begin{cases}-2^{i}\sigma&\text{ if }u=v_{\alpha_{i}}\text{ and }w\in V(G)\\ 2^{i}\sigma&\text{ if }u=v_{\beta_{i}}\text{ and }w\in V(G)\\ 2^{i+j}\sigma&\text{ if }u=v_{\alpha_{i}}\text{ and }w=v_{\alpha_{j}}\\ 2^{i+j}\sigma&\text{ if }u=v_{\beta_{i}}\text{ and }w=v_{\beta_{j}}\\ \end{cases}$
For $u\in V(G)$ and $w\in V(G)$, we have
$c_{uw}=\begin{cases}2\sigma-\gamma_{d}&\text{ if }uw\in E(G)\\ 2\sigma&\text{ if }uw\not\in E(G)\end{cases}$
Edges not specified above have weight zero. Let the penalty parameter be $\sigma=\gamma n+1$. Then all weights are integers and it holds that $Q(\gamma)=\mathtt{offset}-\operatorname{max-cut}(G^{\prime\prime})$ where
[…]
6 Numerical results
All of our algorithms were written [^1] The code is available on the arXiv page of this paper and on https://github.com/melaniesi/EdgeExpansion.jl in Julia [6] version 1.9.2. That is, the split-and-bound Algorithm 2 including pre-elimination and the transformation from $k$-bisection to max-cut problems. Also, Algorithm 3 we implemented in Julia. The SDPs to compute our cheap lower bounds $\ell_{k}$ from the bisection problem in 7 are solved with MOSEK 10.0 [30] using JuMP [26]. We also use JuMP to solve MIQCPs with Gurobi [15] version 11.0. The solver BiqBin [16] for binary quadratic problems was used to solve the parametrized problems in Dinkelbach’s method, and we extended the C code of this solver by adding the option to provide an initial lower bound on the maximization problem. The corresponding changes are tracked in the git repository https://gitlab.aau.at/BiqBin/biqbin. All computations were carried out on an AMD EPYC 7532 with 32 cores with 3.30GHz and 1024GB RAM, operated under Debian GNU/Linux 11.
6.1 Benchmark instances
Randomly generated 0/1-polytopes
The first class of graphs are the graphs of random 0/1-polytopes. The polytopes are generated by randomly selecting $n_{d}$ vertices of the polytope in dimension $d$, i.e., $n_{d}$ different 0/1-vectors in dimension $d$. To obtain the graph, we then solve a linear programming feasibility problem to check whether there is an edge for a given pair of vertices. For any pair $(d,n_{d})$ with $d\in\{8,9,10\}$ and $n_{8}\in\{164,189\}$, $n_{9}\in\{153,178,203,228,253,278\}$, and $n_{10}\in\{256,281\}$, we generated 3 random 0/1-polytopes. The choice of these values is motivated by the aim to generate graphs with 150 to 300 vertices, and to randomly sample between 25% and 75% of the possible 0/1-vectors, with percentages chosen in incremental steps within this range.
Grlex and grevlex graphs
Another class of graphs we consider are the graphs of grlex and grevlex polytopes introduced and characterised by Gupte and Poznanović [13]. The grevlex-$d$ and grlex-$d$ instances of our benchmark set are the corresponding graphs of the polytopes in dimension $d$. It was shown in [13] that those graphs have a very specific structure and that the edge expansion of all grlex-$d$ instances is 1.
DIMACS and Network graphs
The last category of graphs originates from the graph partitioning and clustering application. While the previous benchmark instances are graphs of polytopes, these instances model relations and networks. The set of DIMACS instances are the graphs of the 10th DIMACS challenge on graph partitioning and graph clustering [4] with at most 500 vertices. Additionally, we consider some more network graphs obtained from the online network repository [33]. We in particular chose connected graphs with structural properties such as multiple clusters or having some vertices of high degree and several vertices with small degree.
6.2 Discussion of the experiments
We compare different algorithms for computing the edge expansion of a graph, namely
  1. 1.
    Split & bound Algorithm 2,
  2. 2.
    Fractional programming using Discrete Newton-Dinkelbach’s method in Algorithm 3,
  3. 3.
    Gurobi for solving the MIQCP.
Algorithm 2 vs. Algorithm 3 vs. Gurobi
The detailed results of our experiments are given in § 6.2, 6.2, 6.2, 6.2 and 6.2. In each of the tables, the first column gives the name of the instance followed by the number of vertices and edges. Column 4 reports the optimal solution, i.e., the edge expansion of the graph.
In the split & bound section of the table, the first two columns give the global lower and upper bound after the pre-elimination Algorithm 1. The number of candidates for $k$ after the pre-elimination is given in column 3. In column 4 we report the number of indices $k\in\mathcal{I}$ we were able to eliminate after solving the root node of the branch-and-bound tree. Column 5 lists the total number of branch-and-bound nodes in the max-cut algorithm for all values of $k$ considered. The last two columns display the time spent in the pre-elimination and the total time (including pre-elimination) of the algorithm.
In the section for Dinkelbach’s algorithm, the first column gives the first guess for the edge expansion, i.e., the first trial for $\gamma$. As described before, we take the upper bound from the heuristic for this initialization. Note, that this first guess may differ from $u^{*}$, since in the pre-elimination step of split & bound we perform 30 additional rounds of simulated annealing for all indices $k\in\mathcal{I}$. Column 2 indicates how many parametrized problems $P_{\gamma_{i}}$ have been solved, and column 3 gives the total number of branch-and-bound nodes for solving all parametrized problems. The fourth column of the results of Dinkelbach’s algorithm displays the total time, including running the heuristic to obtain the first guess.
The final column of the tables holds information about computing the edge expansion using Gurobi. For the graphs from the randomly generated polytopes, Gurobi did not succeed to solve any of these instances within a time limit of 3 hours, we therefore report the gap after this time limit in § 6.2. In § 6.2, 6.2, 6.2 and 6.2 we report the time for computing the edge expansion.
We highlight in the tables, which of our two algorithms performs better.
Algorithm 2 (Split & bound)
As can be seen in all tables, the pre-elimination phase of split & bound only leaves a comparably small number of candidates for $k$ to be further investigated. Remember that the number of potential candidates is $\lfloor\frac{n}{2}\rfloor$, whereas $|\mathcal{I}|$ is the number of candidates that remain after the pre-elimination. For 12 instances we were able to compute the edge expansion even within the pre-elimination phase, i.e. the set of candidates $\mathcal{I}$ was empty. For the randomly generated 0/1-polytopes on average only 12% of the candidates have to be further examined, for the other instance classes we can approximately eliminate 80% of the candidate values for $k$ on average. Only for 6 instances we were not able to half the number of candidates within the pre-elimination phase. All of those 6 instances are from the DIMACS or network graphs test set. This indicates that in general already the cheap SDP bound is of good quality.
We also observe that the SDP bound in the root-node of the branch-and-bound tree is of high quality: in 48 out of the 67 instances the gap is closed within the root node for all candidates to be considered. For the other instances the average percentage of candidates left after the root is 8%, where only for one instance the number of remaining candidates is still above 20%.
The heuristic for computing upper bounds also performs extremely well: for almost all instances the upper bound found is the edge expansion of the graph, see columns titled $h(G)$ and $u^{*}$. In fact, only for 3 instances the heuristic fails to find the optimal solution.
Overall, we solve almost all of the considered instances within a few minutes, for very few instances the branch-and-bound tree grows rather large and therefore computation times exceed several hours.
Algorithm 3 (Discrete Newton-Dinkelbach)
Whenever the heuristic already returns the value of the optimal solution, we only have to solve one parametrized problem to certify the optimality of this value. For most of the instances tested, this certificate is already obtained in the root node of the branch-and-bound tree. However, there are many instances where BiqBin terminates because of numerical problems even for the first parametrized problem, see § 6.2. This in particular arises when $\gamma_{d}$ and $\gamma_{n}$ (see Section § 5.1) are large.
Solving the MIQCP with Gurobi
To compute the edge expansion using Gurobi, we input the last formulation of (1) adding the redundant constraint $y\geq 0$. Without this constraint, Gurobi terminated only after 1.65 hours/3 548 work units (resp. more than 24 hours/59 000 work units) on a graph with 29 vertices and 119 edges (resp. 37 vertices and 176 edges) corresponding to the grevlex polytope in dimension 7 (resp. 8).
Adding the redundant constraint, Gurobi is very efficient for graphs with an expansion less than one, see § 6.2 and 6.2, but as soon as the expansion (and also the number of vertices of the graph) gets larger, Gurobi cannot solve the instance within a few hours, see § 6.2 and 6.2.
Impact of the edge expansion
The performance of an algorithm is not in particular depending on their size or density. As noted above, Gurobi is very efficient on graphs with an edge expansion less than one. As the expansion gets larger, our algorithms are the clear winners over Gurobi.
As for the performance of Algorithm 3, we observe that for large expansion the algorithm performs very well in general. But we observe, that in particular if the edge expansion is a fraction of large (coprime) nominator and denominator, this is a disadvantage of the algorithm.
For split & bound, there seems to be no impact of the edge expansion on the performance.
Conclusion
To summarize the results, we give a performance profile in Figure 3. Gurobi solves the MIQCP very efficiently for several instances, but fails to yield results for others within a time limit of 3 hours. It is the clear winner for instances with very small edge expansion. Comparing split & bound with the algorithm following the Discrete Newton-Dinkelbach method, we observe the following behavior. For the grlex instances, Dinkelbach performs extremely well compared to the split & bound approach, see § 6.2. Whereas for the grevlex instances in § 6.2, we observe that, except for dimension 13, the split & bound algorithm by far outperforms Algorithm 3. In addition to the already mentioned, there are some other instances where the difference in the runtimes between the two algorithms is significant. For example, on the instances rand01-9-153-0 and malaria_genes_HVR1 split & bound clearly dominates Algorithm 3, whereas the latter is significantly better on the instances rand01-9-2781 and sp-office.
The conclusion is that in general for graphs with larger edge expansion, the split & bound algorithm is best, and for graphs with small edge expansion Algorithm 3 has a better performance than split & bound, but there are a few exceptions, and the difference in the total time of solving an instance can be quite large.
Overall, we conclude that with our algorithms we can compute the edge expansion of various graphs of size up to around 400 vertices and no other algorithm can achieve this. The time for solving an instance varies, it can be a few seconds for very structured instances and it can exceed one hour, in particular for the instances coming from 0/1-polytopes with rather large expansion. For standard branch-and-cut solvers like Gurobi these instances are out of reach.
Algorithm 2 (split & bound) Algorithm 3 (Dinkelbach) Gurobi
Instance
$n$ $m$ $h(G)$ $\min\,\ell_{k}$ $u^{*}$ $\lvert\mathcal{I}\rvert$
solved
in root
B&B
nodes
Alg. 1
time (s)
total
time (s)
first
guess
# of
steps
B&B
nodes
total
time (s)
relative
gap
[…]
Table 2: Comparison of Algorithm 2 (split & bound), Algorithm 3 (Dinkelbach) and Gurobi for graphs of random 0/1-polytopes. The last column displays the gap reported by Gurobi after a time limit of 3 hours. * Increased timelimit of Gurobi to match the time one of the other algorithms needed for exact computation.
Algorithm 2 (split & bound) Algorithm 3 (Dinkelbach) Gurobi
Instance
$n$ $m$ $h(G)$ $\min\,\ell_{k}$ $u^{*}$ $\lvert\mathcal{I}\rvert$
solved
in root
B&B
nodes
Alg. 1
time (s)
total
time (s)
first
guess
# of
steps
B&B
nodes
total
time (s)
total
time (s)
[…]
Table 3: Comparison of Algorithm 2 (split & bound) and Algorithm 3 (Dinkelbach) for grlex instances.
Algorithm 2 (split & bound) Algorithm 3 (Dinkelbach) Gurobi
Instance
$n$ $m$ $h(G)$ $\min\,\ell_{k}$ $u^{*}$ $\lvert\mathcal{I}\rvert$
solved
in root
B&B
nodes
Alg. 1
time (s)
total
time (s)
first
guess
# of
steps
B&B
nodes
total
time (s)
total
time (s)
[…]
Table 4: Comparison of Algorithm 2 (split & bound), Algorithm 3 (Dinkelbach) and Gurobi for grevlex instances.
Algorithm 2 (split & bound) Algorithm 3 (Dinkelbach) Gurobi
Instance
$n$ $m$ $h(G)$ $\min\,\ell_{k}$ $u^{*}$ $\lvert\mathcal{I}\rvert$
solved
in root
B&B
nodes
Alg. 1
time (s)
total
time (s)
first
guess
# of
steps
B&B
nodes
total
time (s)
total
time (s)
[…]
Table 5: Comparison of Algorithm 2 (split & bound), Algorithm 3 (Dinkelbach) and Gurobi for DIMACS instances.
Algorithm 2 (split & bound) Algorithm 3 (Dinkelbach) Gurobi
Instance
$n$ $m$ $h(G)$ $\min\,\ell_{k}$ $u^{*}$ $\lvert\mathcal{I}\rvert$
solved
in root
B&B
nodes
Alg. 1
time (s)
total
time (s)
first
guess
# of
steps
B&B
nodes
total
time (s)
total
time (s)
[…]
Table 6: Comparison of Algorithm 2 (split & bound), Algorithm 3 (Dinkelbach) and Gurobi for network instances.
00100100100100200200200200300300300300400400400400500500500500000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111time (sec)of instances solvedGurobiDinkelbachSplit & Bound
[…]
Figure 3: Performance comparison of the exact algorithms. Note the different scale on the $x$-axis: the plot on the left displays the time range from 0 to 500 seconds, the plot on the right from 500 seconds to 3 hours.
7 Summary and future research
We developed a split & bound algorithm as well as an algorithm applying Dinkelbach’s idea for fractional programming to compute the edge expansion of a graph. The splitting refers to the fact, that we consider the different values of $k$ ($k$ being the size of the smaller partition) separately. We used semidefinite programming in both phases of our algorithm: on the one hand, SDP-based bounds are used to eliminate several values for $k$ and we use an SDP-based max-cut solver to solve the problem for $k$ fixed. Also, the algorithm following the Dinkelbach framework uses semidefinite programming in order to solve the underlying parametrized problems. Through numerical results on various graph classes, we demonstrate that our split-and-bound algorithm is a robust method for computing the edge expansion while Dinkelbach’s algorithm and Gurobi are very sensitive with respect to the edge expansion of the graph.
In some applications, one is not interested in the exact value of the edge expansion but wants to check whether a certain value is a lower bound on the edge expansion, e.g., to check the Mihail-Vazirani conjecture on 0/1-polytopes. We implemented an option in our algorithm that enables this feature of verifying a given lower bound.
As a heuristic, we use a simulated annealing approach that works very well for the problem sizes we are interested in. However, if one wants to obtain high-quality solutions for larger instances, a more sophisticated heuristic will be needed. Tabu-search, genetic algorithms, or a heuristic in the spirit of the Goemans-Williamson rounding could be potential candidates. In future research, we will also investigate convexification techniques by using recent results on fractional programming [18] and on exploiting submodularity of the cut function as has been done for mixed-binary conic optimization [3].
Appendix A Computing the maximum cut of a graph
Some of our algorithms for computing $h(G)$ rely on finding the maximum cut in a graph. For computing the value of the max-cut, we will use the SDP-based solver BiqBin [16]. Note that the software BiqBin can not only compute the max-cut in a graph and solve quadratic unconstrained binary problems (QUBOs) but it is also applicable to linearly constrained binary problems with a quadratic objective function. However, we only need the max-cut solver in this work, and briefly describe the main ingredients in this section.
BiqBin is a branch-and-bound algorithm that uses a tight SDP relaxation as upper bound and the celebrated Goemans-Williamson rounding procedure to generate a high-quality lower bound on the value of the maximum cut in a graph. To be more precise, the SDP
$\max_{X}\,\Big{\{}\frac{1}{4}\langle L,X\rangle:\operatorname{diag}(X)=e,\,\mathcal{A}(X)=b,\,X\succeq 0\Big{\}}$ (12)
where $\mathcal{A}(X)\leq b$ models a set of triangle-, pentagonal- and heptagonal-inequalities is approximately solved using a bundle method. To do so, only the inequality constraints are dualized yielding the nonsmooth convex partial dual function
$f(\gamma)$ $=\max_{X}\Big{\{}\frac{1}{4}\langle L,X\rangle-\gamma^{\top}(\mathcal{A}(X)-b):\operatorname{diag}(X)\ =\ e,\,X\succeq 0\Big{\}}$
$=b^{\top}\gamma+\max_{X}\Big{\{}\Big{\langle}\frac{1}{4}L-\mathcal{A}^{\top}(\gamma),X\Big{\rangle}:\operatorname{diag}(X)=e,\,X\succeq 0\Big{\}}$
where $\gamma$ are the nonnegative dual variables associated with the constraints $\mathcal{A}(X)\leq b$. Evaluating the dual function $f(\gamma)$ and computing the subgradient amounts to solving an SDP that can be efficiently computed using an interior-point method tailored for this problem. It provides us with the matching pair $(X_{\gamma},\gamma)$ such that $f(\gamma)=b^{\top}\gamma+\langle\nicefrac{{1}}{{4}}L-\mathcal{A}^{\top}(\gamma),X_{\gamma}\rangle$. Moreover, the subgradient of $f$ at $\gamma$ is given by $\partial f(\gamma)=b-\mathcal{A}(X_{\gamma}).$ For obtaining an approximate minimizer of problem
$\min_{\gamma}\{f(\gamma):\gamma\geq 0\},$
the bundle method is used. We refer to [35] for more details.
Note that interior-point methods are far from computing a solution of 12 already for small graphs due to the number of constraints being too large and therefore already forming the system matrix is an expensive task or even impossible due to memory requirements.
BiqBin dominates all max-cut solvers based on linear programming and is comparable to the SDP-based solver BiqCrunch [23]. Moreover, BiqBin is available as a parallelized version.